If xy+yz+zx = 1 Then X/(1+X^2) + Y/(1+Y^2) + Z/(1+Z^2) IS 

Grade:8

1 Answers

Saurabh Koranglekar
5 years ago
Dear student

Consider x=y=z=1/(3)^0.5

X/(1+X^2) + Y/(1+Y^2) + Z/(1+Z^2) = (3*( 1/3^0.5 )) ./ (1+(1/3))

= (3(3)^0.5)/4